x[squared]=9-y[squared]
Answer by stanbon(48579)

x[squared]=9-y[squared]
--------
x^2 = 9 - y^2
So, x^2 + y^2 = 9
The graph is a circle with its center
at (0,0) having a radius of 3.
----
Cheers,
Stan H.